### v1

# Jl. Ganesha 10, Bandung 40132, Indonesia

Abstract
Galaxy clusters are the largest virialized structures in the Universe and are predominantly dominated by dark matter. The hydrostatic mass and the mass obtained from gravitational lensing measurements generally differ, a discrepancy known as the hydrostatic mass bias. In this work, we derive the hydrostatic mass of galaxy clusters within the framework of Rastall gravity and investigate its implications under two scenarios: (i) the absence of dark matter and (ii) the existence of dark matter. In the first scenario, Rastall gravity effectively reduces the hydrostatic mass, bringing it closer to the observed baryonic mass. The best linear fit yields a slope $\mathbf{M}=1.07\pm 0.11$, indicating a near one-to-one correspondence between the two masses. In the second scenario, Rastall gravity helps to alleviate the hydrostatic mass bias. The linear fit between the Rastall hydrostatic mass and the observed lensing mass results in a best-fit slope $\mathbf{M}=1.01\pm 0.16$, which is very close to unity. These results suggest that Rastall gravity provides a statistically favorable framework for addressing mass discrepancies in galaxy clusters.
1 Introduction
Galaxy clusters are the largest virialized astronomical objects in the universe [1]. The formation of galaxy clusters is related to the collapse of the largest gravitationally bound overdensities in the initial density field and is accompanied by the most energetic phenomena since the Big Bang and by the complex interplay between gravity-induced dynamics of collapse and baryonic processes associated with galaxy formation [2]. The population of galaxies in clusters has evolved rapidly over the last 5 Gyr [3]. Galaxy clusters host a hot and diffuse intracluster medium, characterized by typical temperatures $\sim 7\times 10^{7}$ K and electron density $\sim 10^{-3}$ cm-3 [4, 5].
The mass of galaxy clusters can be determined using three main methods: velocity dispersion measurements, X-ray or Sunyaev-Zel’dovich (SZ) observations under the assumption of hydrostatic equilibrium, and gravitational lensing. Except for gravitational lensing, these methods estimate the cluster mass within the framework of the Newtonian approximation [4, 6]. The reviews of the aforementioned methods are available in Refs. [2, 7, 8, 9, 10, 11, 12, 13, 14, 15].
It should be noted that the lensing mass of galaxy clusters is systematically higher than the hydrostatic mass. This discrepancy is commonly referred to as the hydrostatic mass bias [16]. In terms of mass composition, galaxy clusters are dominated by dark matter (DM) [17], which can account for approximately $80-90\%$ of their total mass [18]. Based on computational simulations, galaxy clusters are identified as the largest gravitationally bound DM halos in the large-scale structure of the Universe [19]. These facts highlight the challenges in accurately determining the masses of galaxy clusters.
Bambi [20] conducted a detailed study on the mass determination of galaxy clusters. Assuming that dark energy (DE) behaves as a cosmological constant $\Lambda$ on relatively small scales, he investigated the effects of a nonzero cosmological constant in the Newtonian limit. As a result, he derived a general relativity (GR) correction to the mass of galaxy clusters, leading to an effective mass for galaxy clusters. Desai [16] computed the masses of several galaxy clusters using Bambi’s formulation and showed that, within the considered samples, the impact of general relativistic corrections on the hydrostatic masses of galaxy clusters is negligible.
Gupta and Desai [6] also evaluated the general relativistic correction to the mass of galaxy clusters using the Tolman-Oppenheimer-Volkoff (TOV) equation, which represents hydrostatic equilibrium in general relativity. They similarly found that the GR correction is negligible, indicating that the standard Newtonian approach provides an adequate description.
On the other hand, the domination of the DE and DM in the our dark Universe has become a motivation to construct alternative theories beyond GR [21]. Rahvar and Mashhoon [22] calculated the masses of galaxy clusters within the framework of nonlocal gravity. They performed the linear fit between the nonlocal and baryonic mass estimates. The ideal scenario, in which nonlocal gravity fully accounts for the missing mass without involving DM within the scale of clusters, corresponds to $\mathbf{M}=1$. By comparing the resulting nonlocal masses with the corresponding baryonic masses for a sample of ten galaxy clusters, the best-fit slope was found to be $\mathbf{M}=0.84\pm 0.04$. Using a similar approach but within the framework of Scalar–Tensor–Vector Gravity (STVG), Moffat and Rahvar [23] obtained a value of $\mathbf{M}$ that is closer to unity than that reported by Rahvar and Mashhoon. They conclude that STVG is compatible with observational data on megaparsec scales without invoking DM.
Another attempt to utilize modified theories of gravity to calculate the masses of galaxy clusters was carried out by the authors in Ref. [4]. They employed Eddington-inspired Born–Infeld (EiBI) theory, beyond Horndeski gravity (BHG), and Newtonian gravity modified by the generalized uncertainty principle (GUP). Among these theories, EiBI theory yields the best linear fit with $\mathbf{M}=0.126\pm 0.086$. This value is significantly lower than that reported by Rahvar and Mashhoon and is also much farther from unity. Consequently, EiBI theory, BHG, and Newtonian gravity modified by the GUP cannot satisfactorily account for the masses of galaxy clusters without involving DM.
Another modified theory of gravity that is worth considering is Rastall gravity, which was introduced by Rastall [24] in 1972. It allows for a nonconserved energy-momentum tensor (EMT), characterized by a nonvanishing covariant divergence of the EMT, which Rastall proposed as a generalization of GR. The violation of EMT conservation can be phenomenologically caused by quantum effects in classical context [25, 26, 27]. In contrast, the covariant divergence of the EMT identically vanishes in GR.
From a different perspective, Visser [28] claimed that Rastall gravity is equivalent to GR. He argued that, in the matter sector, Rastall’s EMT can be interpreted as an artificially isolated component of a physically conserved EMT. Consequently, the resulting theory does not introduce new gravitational dynamics and is therefore equivalent to GR.
Darabi et al. [29] refuted Visser’s viewpoint. They emphasized that Rastall did not introduce new forms of the EMT in his original work. Instead, Rastall proposed a modified relation between the EMT and spacetime geometry, motivated by a possible unknown mutual interaction between them. They further stressed that violations of energy-momentum conservation have also been considered in theoretical frameworks distinct from Rastall gravity, for instance in attempts to model DE.
Despite this debate, Rastall gravity has been widely explored in the literature on modified theories of gravity. It has been applied to various astrophysical objects, including neutron stars [25, 27, 30, 31, 32], quark stars [33, 34, 35, 36, 37], wormholes [38, 39, 40, 41, 42, 43], black holes [26, 44, 45, 46, 47, 48, 49, 50, 51], and gravastars [52], as well as to cosmology in the context of an accelerated expanding Universe [53]. On galactic scales, Rastall gravity has been investigated by Li et al. [54], who modeled both stellar matter and DM as perfect fluids. Tang et al. [55] constrained the Rastall parameter using rotation curves of low surface brightness galaxies.
The aims of this work are twofold. First, we construct the formalism for the hydrostatic mass of galaxy clusters within the framework of Rastall gravity and determine the best-fit value of $\mathbf{M}$, following the methodology adopted in Refs. [4, 22, 23], with the goal of assessing whether a viable mass modeling can be achieved without invoking dark matter. Second, we determine the best-fit value of $\mathbf{M}$ in order to investigate whether the hydrostatic mass bias can be alleviated within this framework, an issue that was not resolved in Ref. [16]. We find that the Rastall parameter has a nontrivial influence on the hydrostatic mass of galaxy clusters, and that the resulting best-fit values of $\mathbf{M}$ are close to unity, with $\mathbf{M}=1$ lying within the associated uncertainty range.
The structure of this paper is as follows. Section 2 presents the mathematical formalisms employed in this work, including Rastall gravity, the standard hydrostatic formulation for determining the mass of galaxy clusters, and the hydrostatic mass formulation within the framework of Rastall gravity. Section 3 presents the numerical results, followed by a discussion of the corresponding implications. Finally, our conclusions are summarized in Section 4.
2 Formalisms
To ensure completeness, we present the formalism of Rastall gravity in Subsection 2.1, the standard hydrostatic formulation for determining the mass of galaxy clusters in Subsection 2.2, and the formulation of the hydrostatic mass of galaxy clusters within the framework of Rastall gravity derived in Subsection 2.3.
2.1 Rastall gravity
First, in this Subsection, we use $G=c=1$, where $G$ is the universal gravitational constant, while $c$ denotes the speed of light. Also, we emphasize that, unlike most modified theories of gravity, which are formulated starting from an action principle, Rastall gravity is instead based on a modification of the divergence of the EMT. In GR, it is widely known that the divergence of EMT $T^{\mu\nu}$ reads
$\nabla_{\mu}T^{\mu\nu}=0,$ (1)
which ensures the EMT conservation. However, Rastall [24] proposed a modification to the divergence of the EMT, given by
$\nabla_{\mu}T^{\mu\nu}=\lambda\nabla^{\nu}\mathcal{R},$ (2)
where $\lambda$ denotes the Rastall free parameter and $\mathcal{R}$ is the Ricci scalar. It is evident that in the limit $\lambda=0$, Rastall gravity is reduced to GR. Thus, $\lambda$ presents the deviation of Rastall gravity from GR.
With Eq. (2) in our hands, we can obtain the Rastall field equation, which writes [27]
$G^{\mu}_{\nu}+\kappa\lambda\mathcal{R}\delta^{\mu}_{\nu}=\kappa T^{\mu}_{\nu}.$ (3)
Here, $G^{\mu}_{\nu}=R^{\mu}_{\nu}-\frac{1}{2}\mathcal{R}\delta^{\mu}_{\nu}$ denotes the Einstein tensor, $\mathcal{R}^{\mu}_{\nu}$ denotes the Ricci tensor, $\delta^{\mu}_{\nu}$ is the Kronecker delta, and $\kappa$ is the constant of the field equation. For 4-dimensional spacetime, Eq. (3) gives
$\mathcal{R}=\frac{\kappa T}{4\kappa\lambda-1},$ (4)
where $T$ denotes the trace of EMT. Consequently, the Rastall field equation now becomes
$G^{\mu}_{\nu}=\kappa\mathcal{T}^{\mu}_{\nu},$ (5)
where $\mathcal{T}^{\mu}_{\nu}$ is the effective modified EMT that satisfies
$\mathcal{T}^{\mu}_{\nu}=T^{\mu}_{\nu}-\delta^{\mu}_{\nu}\left(\frac{\kappa\lambda}{4\kappa\lambda-1}T\right).$ (6)
Before proceeding, it is worth pointing out that Refs. [25, 26, 46] adopted $\kappa=8\pi$ in Rastall gravity, as in GR. Nevertheless, the constant in Rastall gravity is intrinsically modified and differs from its GR counterpart. The authors of Ref. [31] derived the constant $\kappa$ in Rastall gravity by applying the theory in the weak-field limit. They obtained
$\kappa=\frac{8\pi}{2\gamma+1},$ (7)
where the parameter $\gamma$ is related to the Rastall parameter $\lambda$ through
$\lambda=\frac{\gamma}{\kappa(4\gamma-1)}.$ (8)
From Eqs. (4), (7), and (8), we have an insight that $\kappa\lambda=1/4$, $\gamma=-1/2$, and $\kappa\gamma=1/4$ are forbidden.
2.2 Standard hydrostatic mass calculation
We start from the hydrostatic equilibrium equation, given by [56]
$\frac{dp}{dr}=-\rho\frac{Gm(r)}{r^{2}},$ (9)
where $p$ is the pressure, $\rho$ denotes the density of the distribution, $m(r)$ is the mass of the clusters enclosed within a radius $r$. Here, the equation of state (EOS) of the galaxy clusters is assumed to be satisfied by the EOS of the ideal gas. Such EOS is employed by Ref. [4, 6]. For the ideal gases, we have [4]
$\frac{dp}{dr}=\frac{k_{B}}{m_{p}\mu}\left(\tau\frac{d\rho}{dr}+\rho\frac{d\tau}{dr}\right).$ (10)
Here, $k_{B}$ denotes the Boltzmann constant, $m_{p}$ is the mass of proton, $\mu$ denotes the mean molecular weight of the cluster, $\tau$ is the absolute temperature.
By combining Eq. (9) and Eq. (10), one obtains
$-\rho\frac{Gm(r)}{r^{2}}$ $=$ $\frac{k_{B}}{m_{p}\mu}\left(\tau\frac{d\rho}{dr}+\rho\frac{d\tau}{dr}\right)$
$m(r)$ $=$ $-\frac{k_{B}r^{2}}{m_{p}\mu G\rho}\left(\tau\frac{d\rho}{dr}+\rho\frac{d\tau}{dr}\right)$
$m(r)$ $=$ $-\frac{k_{B}r\tau}{m_{p}\mu G}\left(\frac{r}{\rho}\frac{d\rho}{dr}+\frac{r}{\tau}\frac{d\tau}{dr}\right).$ (11)
It is widely known that
$\left(\frac{r}{\rho}\frac{d\rho}{dr}+\frac{r}{\tau}\frac{d\tau}{dr}\right)=\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right).$ (12)
By substituting Eq. (12) into Eq. (11), one can obtain
$m(r)=-\frac{k_{B}r}{\mu m_{p}G}\tau\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right).$ (13)
By substituting the values of the constants $k_{B}$, $\mu$, $m_{p}$, and $G$, we have the expression of the hydrostatic mass of galaxy clusters, i.e.
$m(r)=-3.7\times 10^{-4}\tau\>r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right)\mathcal{M}_{\odot}.$ (14)
Here, $\mathcal{M}_{\odot}=10^{14}\>M_{\odot}$.
Following Ref. [22], the hot gas density of galaxy clusters is given by
$\rho\approx 1.24\,m_{p}\,[n_{e}(r)n_{p}(r)]^{1/2},$ (15)
where $n_{e}$ and $n_{p}$ denote the number densities of electrons and protons, respectively. The product $n_{e}(r)n_{p}(r)$ can be expressed as
$n_{e}(r)n_{p}(r)=\frac{(r/r_{c})^{-\alpha^{\prime}}}{(1+r^{2}/r_{r}^{2})^{3\beta-\alpha^{\prime}/2}}\frac{n_{0}^{2}}{(1+r^{\gamma}/r_{s}^{\gamma})^{\varepsilon/\gamma}}+\frac{n_{0}^{\prime 2}}{(1+r^{2}/r_{c}^{\prime 2})^{2\beta^{\prime}}}.$ (16)
Table 1 represents the parameter related to the density for ten galaxy clusters.
Table 1: Parameters for the gas density profiles of the galaxy clusters. The data are taken from Ref. [22].
Cluster $r_{500}$ $n_{0}$ $r_{c}$ $r_{s}$ $\alpha^{\prime}$ $\beta$ $\varepsilon$ $n_{0}^{\prime}$ $r_{c}^{\prime}$ $\beta^{\prime}$
(kpc) ($10^{-3}\,\mathrm{cm^{-3}}$) (kpc) (kpc) ($10^{-1}\,\mathrm{cm^{-3}}$) (kpc)
A133 $1007\pm 41$ 4.705 94.6 1239.9 0.916 0.526 4.943 0.247 75.83 3.607
A383 $944\pm 32$ 7.226 112.1 408.7 2.013 0.577 0.767 0.002 11.54 1.000
A478 $1337\pm 58$ 10.170 155.5 2928.9 1.254 0.704 5.000 0.762 23.84 1.000
A907 $1096\pm 30$ 6.257 136.9 1887.1 1.556 0.594 4.998
A1413 $1299\pm 43$ 5.239 195.0 2153.7 1.247 0.661 5.000
A1795 $1235\pm 36$ 31.175 38.2 682.5 0.195 0.491 2.606 5.695 3.00 1.000
A1991 $732\pm 33$ 6.405 59.9 1064.7 1.735 0.515 5.000 0.007 5.00 0.517
A2029 $1362\pm 43$ 15.721 84.2 908.9 1.164 0.545 1.669 3.510 5.00 1.000
A2390 $1416\pm 48$ 3.605 308.2 1200.0 1.891 0.658 0.563
MKW4 $634\pm 28$ 0.196 578.5 595.1 1.895 1.119 1.602 0.108 30.11 1.971
According to Ref. [22], the intracluster plasma can be divided into two regions: a cooling zone near the cluster center and an outer region. The temperature profiles in these two regions are modeled by two different functions. At the center of a cluster the temperature decreases due perhaps to radiative cooling in this region; hence
$\tau_{in}(r)=\frac{x_{0}+T_{min}/T_{0}}{x_{0}+1},$ (17)
where
$x_{0}=\left(\frac{r}{r_{cool}}\right)^{a_{cool}}.$ (18)
For the outside region of the central coolong zone,
$\tau_{out}(r)=\frac{(r/r_{t})^{a^{\prime}}}{\left[1+\left(\frac{r}{r_{t}}\right)^{b}\right]^{c^{\prime}/b}}.$ (19)
Here, $r_{t}$ denotes the radial transision region. Temperature profile of the cluster reads
$\tau(r)=T_{0}\tau_{in}(r)\tau_{out}(r).$ (20)
The parameters related to the temperature profile are given in Table 2
Table 2: The parameters of the temperature profiles for ten galaxy clusters. The data are taken from Ref. [22].
Cluster $T_{0}$ $r_{t}$ $\alpha^{\prime}$ $b$ $c^{\prime}$ $T_{\min}/T_{0}$ $r_{\mathrm{cool}}$ $\alpha_{\mathrm{cool}}$
(keV) (Mpc) (kpc)
A133 3.61 1.42 0.12 5.00 10.0 0.27 57 3.88
A383 8.78 3.03 -0.14 1.44 8.0 0.75 81 6.17
A478 11.06 0.27 0.02 5.00 0.4 0.38 129 1.60
A907 10.19 0.24 0.16 5.00 0.4 0.32 208 1.48
A1413 7.58 1.84 0.08 4.68 10.0 0.23 30 0.75
A1795 9.68 0.55 0.00 1.63 0.9 0.10 77 1.03
A1991 2.83 0.86 0.04 2.87 4.7 0.48 42 2.12
A2029 16.19 3.04 -0.03 1.57 5.9 0.10 93 0.48
A2390 19.34 2.46 -0.10 5.00 10.0 0.12 214 0.08
MKW4 2.26 0.10 -0.07 5.00 0.5 0.85 16 9.62
The entity $\frac{d\ln\rho}{d\ln r}$ in Eq. (14) is given by
$\frac{d\ln\rho(r)}{d\ln r}={}$ $\frac{1}{H(r)}\Bigg[-n_{0}^{2}\alpha^{\prime}\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}}$ (21)
$\times\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{0.5\alpha^{\prime}-3\beta^{\prime}}\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-\frac{\varepsilon}{\gamma}}$
$+n_{0}^{2}(\alpha^{\prime}-6\beta)\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}+2}$
$\times\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{-1+0.5\alpha^{\prime}-3\beta}\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-\frac{\varepsilon}{\gamma}}$
$-6n_{0}^{\prime 2}\beta^{\prime}\left(\frac{r}{r_{c}^{\prime}}\right)^{2}\left(1+\frac{r^{2}}{r_{c}^{\prime 2}}\right)^{-1-3\beta^{\prime}}$
$-n_{0}^{2}\varepsilon\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}}\left(\frac{r}{r_{s}}\right)^{\gamma}$
$\times\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{0.5\alpha^{\prime}-3\beta}\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-1-\frac{\varepsilon}{\gamma}}\Bigg].$
where
$H(r)={}$ $2\Bigg[n_{0}^{\prime 2}\left(1+\frac{r^{2}}{r_{c}^{\prime 2}}\right)^{-3\beta^{\prime}}$ (22)
$\quad+n_{0}^{2}\left(\frac{r}{r_{c}}\right)^{-\alpha^{\prime}}\left(1+\frac{r^{2}}{r_{c}^{2}}\right)^{0.5\alpha^{\prime}-3\beta}$
$\quad\times\left(1+r^{\gamma}r_{s}^{-\gamma}\right)^{-\frac{\varepsilon}{\gamma}}\Bigg].$
2.3 Hydrostatic mass of galaxy clusters in Rastall gravity
Rastall gravity is intrinsically formulated within a general relativistic framework, whereas the hydrostatic mass of galaxy clusters is commonly derived in the Newtonian limit. To bridge these two regimes, we first formulate the hydrostatic equilibrium condition in the general relativistic context, i.e. the TOV equation modified by Rastall gravity, and subsequently derive its weak-field limit. Throughout the general relativistic regime formulation, we adopt natural units by setting $G=c=1$. The physical constants will be reinstated when taking the Newtonian limit.
Consider an interior Schwarzschild metric of the form
$ds^{2}=-e^{2\nu(r)}dt^{2}+e^{2\lambda(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta,d\phi^{2},$ (23)
where $e^{2\nu(r)}=1-\dfrac{2m(r)}{r}$. We also consider the EMT of a perfect fluid,
$T^{\mu}_{\ \nu}=\begin{pmatrix}-\rho&0&0&0\\ 0&p&0&0\\ 0&0&p&0\\ 0&0&0&p\end{pmatrix}$. (24)
By applying the Rastall field equations to the metric and the energy–momentum tensor given above, one obtains the Rastall-modified TOV equation, i.e. [27]
$\frac{dp}{dr}=-\frac{\frac{(\rho+p)\left(m+\frac{\kappa}{2}r^{3}\mathcal{P}\right)}{r(r-2m)}}{1-\frac{\kappa\lambda}{4\kappa\lambda-1}\left(3-\frac{d\rho}{dp}\right)},$ (25)
where $\mathcal{P}=p-\left[\frac{\kappa\lambda}{4\kappa\lambda-1}(3p-\rho)\right]$.
Recall that, in GR, the hydrostatic equilibrium equation in the Newtonian regime can be recovered from the TOV equation by assuming $\rho\gg p$ and $r\gg m$. By adopting the same procedure and also considering $\frac{d\rho}{dp}=\frac{\rho}{r}\frac{d\ln\rho}{d\ln r}\left(\frac{dp}{dr}\right)^{-1}$, we obtain the Newtonian-limit expression of Eq. (25), i.e.
$\frac{dp}{dr}=-\frac{\rho}{r^{2}}\left(\frac{m+\frac{1}{2}\frac{\kappa^{2}\lambda}{4\kappa\lambda-1}r^{3}\rho}{1-\frac{\kappa\lambda}{4\kappa\lambda-1}\left[3-\frac{\rho}{r}\frac{d\ln\rho}{d\ln r}\left(\frac{dp}{dr}\right)^{-1}\right]}\right).$ (26)
With a little algebra, we find
$\frac{dp}{dr}=\left[-\frac{\rho}{r^{2}}\left\{m+\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{\kappa r^{3}\rho}{2}+r\frac{d\ln\rho}{d\ln r}\right)\right\}\right]\frac{4\kappa\lambda-1}{\kappa\lambda-1}.$ (27)
Recall that for an ideal gas, we have Eq. (10). Substituting this relation into the Rastall-modified Newtonian-limit hydrostatic equilibrium equation and restoring the gravitational constant $G$ and the speed of light $c$ in the expression, we obtain
$m(r)=-\frac{k_{B}}{\mu m_{p}}\frac{\kappa\lambda-1}{4\kappa\lambda-1}\tau(r)r\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln\tau}{d\ln r}\right)-\frac{\kappa\lambda}{4\kappa\lambda-1}\left(\frac{\kappa r^{3}\rho}{2}+\frac{c^{2}r}{G}\frac{d\ln\rho}{d\ln r}\right).$ (28)
Eq. (28) is the Rastall version of the hydrostatic mass of galaxy clusters. When $\lambda=0$, the equation is reduced to Eq. (13).
(a) (a) Linear fit to the relation between the hydrostatic masses according to standard non-Rastall gravity and the observed baryonic masses, (b) the likelihood function for parameter $\mathbf{M}$ in standard non-Rastall gravity, (c) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=9.55\times 10^{-2}$ and the observed baryonic masses, (d) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=9.55\times 10^{-2}$, (e) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.14\times 10^{-1}$ and the observed baryonic masses, and (f) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.14\times 10^{-1}$.
(b) (a) Linear fit to the relation between the hydrostatic masses according to standard non-Rastall gravity and the observed lensing masses, (b) the likelihood function for parameter $\mathbf{M}$ in standard non-Rastall gravity, (c) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.49\times 10^{-3}$ and the observed baryonic masses, (d) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.49\times 10^{-3}$, (e) linear fit to the relation between the hydrostatic masses according to Rastall gravity with $\lambda=1.69\times 10^{-3}$ and the observed baryonic masses, and (f) the likelihood function for parameter $\mathbf{M}$ in Rastall gravity for $\lambda=1.69\times 10^{-3}$.
3 Numerical results and discussion
Table 3: Comparison between the Newtonian hydrostatic mass, the hydrostatic mass in Rastall gravity for two different values of the Rastall parameter $\lambda$, and the observed baryonic mass of galaxy clusters. The Newtonian hydrostatic mass and $r_{500}$ data are taken from Ref. [4], while the baryonic mass data are taken from Ref. [22].
Cluster $r_{500}$ $M_{\mathrm{Newt}}$ $M_{\mathrm{hyd}}^{\mathrm{R}}$ $M_{\mathrm{hyd}}^{\mathrm{R}}$ $M_{\mathrm{bar}}$
(kpc) $(10^{14}M_{\odot})$ $(10^{14}M_{\odot})$ $(10^{14}M_{\odot})$ $(10^{14}M_{\odot})$
$\lambda=9.55\times 10^{-2}$ $\lambda=1.14\times 10^{-1}$
A133 $1007\pm 41$ 4.1 0.41 0.53 $0.302\pm 0.054$
A383 $944\pm 32$ 3.2 0.32 0.42 $0.412\pm 0.059$
A478 $1337\pm 58$ 7.4 0.75 0.97 $0.982\pm 0.204$
A907 $1096\pm 30$ 4.4 0.44 0.58 $0.599\pm 0.072$
A1413 $1299\pm 43$ 9.9 1.00 1.30 $0.839\pm 0.134$
A1795 $1235\pm 36$ 5.8 0.58 0.75 $0.683\pm 0.090$
A1991 $732\pm 33$ 1.4 0.14 0.18 $0.145\pm 0.027$
A2029 $1362\pm 43$ 8.6 0.87 1.13 $1.026\pm 0.146$
A2390 $1416\pm 48$ 13.0 1.29 1.68 $1.526\pm 0.246$
MKW4 $634\pm 28$ 0.75 0.08 0.10 $0.062\pm 0.008$
Table 4: Comparison between the hydrostatic Newtonian mass, hydrostatic mass in Rastall gravity for two values of the Rastall parameter, and the lensing mass of galaxy clusters. The Newtonian mass and $r_{500}$ data are taken from Ref. [4], while the lensing mass data are taken from Refs. [57, 58, 59].
Cluster $r_{500}$ $M_{\mathrm{Newt}}$ $M_{hyd}$ $M_{hyd}$ $M_{\mathrm{len}}$
(kpc) $(10^{14}M_{\odot})$ $(10^{14}M_{\odot})$ $(10^{14}M_{\odot})$ $(10^{14}M_{\odot})$
$\lambda=1.49\times 10^{-3}$ $\lambda=1.69\times 10^{-3}$
A383 $944\pm 32$ 3.20 3.70 3.80 5.03
A478 $1337\pm 58$ 7.40 8.57 8.80 8.67
A1795 $1235\pm 36$ 5.80 6.65 6.83 9.20
A1991 $732\pm 33$ 1.40 1.60 1.64 2.50
A2029 $1362\pm 43$ 8.60 9.98 10.24 10.80
In this work, we compute the masses of galaxy clusters at the radius $r_{500}$. As shown in Tables 3 and 4, the values of $r_{500}$ are subject to observational uncertainties. In our analysis, however, the mass is evaluated at the central (best-fit) value of $r_{500}$, while the propagation of the uncertainty in $r_{500}$ is not taken into account. This approach is also employed in Ref. [4].
We consider two distinct possibilities. The first possibility assumes the absence of DM; in this case, we fit the hydrostatic mass derived within the framework of Rastall gravity to the observed baryonic mass in order to obtain the best linear fit. The ideal situation corresponds to a slope $\mathbf{M}=1$, which indicates a one-to-one correspondence between the hydrostatic and baryonic masses. This approach has been adopted in Ref. [4] using EiBI theory, BHG, and Newtonian gravity modified by the GUP; in Ref. [22] within the context of nonlocal gravity; and in Ref. [23] using STVG.
The second possibility assumes the existence of DM, but aims to alleviate the mass discrepancy between the hydrostatic and the gravitational lensing masses. In this scenario, we perform a linear fit between the hydrostatic mass in Rastall gravity and the observed lensing mass to obtain the best-fit slope. Similarly to the first case, the ideal situation is achieved when the slope $\mathbf{M}=1$. This approach has been employed in Ref. [16], where an effective mass derived from the Kottler metric was used.
Table 3 presents a comparison between the Newtonian hydrostatic mass, the hydrostatic mass in Rastall gravity for two values of the Rastall parameter, namely $\lambda=9.55\times 10^{-2}$ and $\lambda=1.14\times 10^{-1}$, and the observed baryonic mass of galaxy clusters. Here we consider ten galaxy clusters as the sample. From the values listed in the table, it is evident that the Rastall parameter can reduce the hydrostatic mass in such a way that it approaches the observed baryonic mass.
The formula for the relative likelihood is presented in Appendix A. From Fig. 6(a), it is evident that the standard (non-Rastall) hydrostatic mass yields a slope that deviates significantly from unity, namely $\mathbf{M}=8.24\pm 0.86$. Even when the uncertainty is taken into account, the value $\mathbf{M}=1$ lies well outside the allowed range. Therefore, within the context of the assumption that DM is absent, this case is not viable. In contrast, the assumption of no DM is more compatible with the Rastall gravity framework.
For Rastall gravity with $\lambda=9.55\times 10^{-2}$, the resulting slope is significantly improved compared to the non-Rastall case. Although the value $\mathbf{M}=1$ is still not encompassed by the uncertainty of the slope, it is nevertheless included within the corresponding relative likelihood curve.
In the comparison between the hydrostatic mass and the baryonic mass, the value $\lambda=1.14\times 10^{-1}$ yields the best linear fit, with a slope $\mathbf{M}=1.07\pm 0.11$. The slope is very close to unity, and the value $\mathbf{M}=1$ lies within both the uncertainty range and the corresponding relative likelihood curve. This result indicates that, from a statistical standpoint, Rastall gravity provides a more favorable description of the hydrostatic-baryonic mass relation than standard gravity under the assumption of the absence of DM. Another noteworthy result is that the best-fit slope obtained in this work is closer to unity than those reported in some related studies. In Ref. [4], the best-fit slopes were found to be $0.126\pm 0.086$ for EiBI theory and $0.023\pm 0.119$ for BHG. Meanwhile, Ref. [22] reported a slope of $0.84\pm 0.04$ within the framework of nonlocal gravity.
We also examine whether the hydrostatic mass within the framework of Rastall gravity can help to alleviate the hydrostatic mass bias, an issue that was not resolved in Ref. [16]. To this end, we consider a sample of five galaxy clusters. The values of the Newtonian hydrostatic mass, the Rastall hydrostatic mass, and the corresponding lensing mass for these clusters are presented in Table 4. For this test, we adopt $\lambda=1.49\times 10^{-3}$ and $\lambda=1.69\times 10^{-3}$. Both values lead to an increase in the hydrostatic mass.
The linear fits between the standard (non-Rastall) hydrostatic mass and the observed lensing mass, as well as between the Rastall hydrostatic mass and the observed lensing mass, are shown in Fig. 6(b). The standard non-Rastall hydrostatic mass yields a slope value of $\mathbf{M}=0.85\pm 0.13$. Although this value is already close to unity, we further investigate whether the inclusion of Rastall gravity can improve the linear relation and bring the slope closer to the ideal value $\mathbf{M}=1$, thereby alleviating the hydrostatic mass bias.
In the comparison between the hydrostatic mass and the lensing mass, the value $\lambda=1.69\times 10^{-3}$ yields the best linear fit, with a slope $\mathbf{M}=1.01\pm 0.16$. The slope is very close to unity, even closer to $\mathbf{M}=1$ than the best-fit value obtained from the comparison between the hydrostatic mass and the observed baryonic mass. Moreover, the value $\mathbf{M}=1$ lies within the uncertainty range and is also located very close to the peak of the corresponding relative likelihood curve.
This result indicates that, from a statistical standpoint, Rastall gravity is able to significantly alleviate the hydrostatic mass bias. Nevertheless, a test with a larger sample of galaxy clusters and more lensing mass data is necessary, both to further assess the viability of Rastall gravity and to constrain the value of the Rastall parameter $\lambda$ using observational data. Another noteworthy outcome of this study is that it addresses the issue encountered in Ref. [16], where the hydrostatic mass bias could not be alleviated within the considered framework.
4 Conclusion
In this study, we have derived the formalism of the hydrostatic mass of galaxy clusters within the framework of Rastall gravity. We find that Rastall gravity has a nontrivial influence on the hydrostatic mass of galaxy clusters. Two scenarios are considered: (i) the absence of dark matter (DM), in which the Rastall hydrostatic mass is compared with the observed baryonic mass; and (ii) the existence of DM, where the Rastall hydrostatic mass is employed to alleviate the hydrostatic mass bias.
In the comparison between the Rastall hydrostatic mass and the observed baryonic mass, using a sample of ten galaxy clusters, the best-fit result is obtained for $\lambda=1.14\times 10^{-1}$, yielding $\mathbf{M}=1.07\pm 0.11$. In this case, the unity value $\mathbf{M}=1$, which represents the ideal situation where the Rastall hydrostatic mass statistically matches the observed baryonic mass, lies within the uncertainty range and is also located in the high-likelihood region of the corresponding relative likelihood curve.
On the other hand, using five galaxy clusters as the sample, we find that the best-fit value, $\mathbf{M}=1.01\pm 0.16$, is obtained for $\lambda=1.69\times 10^{-3}$. In this case, unity lies very close to the peak of the corresponding relative likelihood curve in the comparison between the Rastall hydrostatic mass and the observed lensing mass. This result indicates that Rastall gravity can effectively alleviate the hydrostatic mass bias problem. Nevertheless, a larger sample of galaxy clusters in the comparison between the Rastall hydrostatic mass and the gravitational lensing mass is required to further assess the viability of Rastall gravity and to constrain the Rastall parameter $\lambda$ using observational data.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author used ChatGPT (OpenAI) in order to improve the language and readability of the manuscript. After using this tool, he reviewed and edited the content as needed and take full responsibility for the content of the paper.
Acknowledgements
MLP is supported by the Indonesia Endowment Fund for Education Agency (LPDP). FA is supported by Physics Study Program, Faculty of Mathematics and Natural Science Education, Universitas Pendidikan Indonesia. We thank the Theoretical High Energy Physics Group at Institut Teknologi Bandung for discussions, insights, and hospitality.
Appendix A Formula of relative likelihood
In this work, the relative likelihood is calculated using the Gaussian form [22]
$\mathcal{L}_{\mathrm{rel}}=\exp\left[-\frac{1}{2}\left(\frac{\mathbf{M}-\mu}{\sigma}\right)^{2}\right],$ (A1)
where $\mathbf{M}$ denotes the ratio $M_{hyd}/M_{bar}$, $\mu$ is the slope obtained from the linear fit, and $\sigma$ is the uncertainty associated with the fit.
Appendix B Statistical data of the linear fits
Table 5: Linear fitting results between Rastall hydrostatic mass and observed baryonic and lensing masses for different values of the Rastall parameter $\lambda$.
Rastall Hydrostatic Mass vs Observed Baryonic Mass
$\lambda$ Intercept Slope RSS Pearson $r$ $R^{2}$ Adj. $R^{2}$
0 $0.44\pm 0.67$ $8.24\pm 0.86$ $10.75$ $0.96$ $0.92$ $0.91$
$9.55\times 10^{-2}$ $0.05\pm 0.07$ $0.82\pm 0.09$ $0.11$ $0.96$ $0.92$ $0.91$
$1.14\times 10^{-1}$ $0.06\pm 0.09$ $1.07\pm 0.11$ $0.19$ $0.96$ $0.92$ $0.91$
Rastall Hydrostatic Mass vs Observed Lensing Mass
0 $0.85\pm 1.01$ $0.85\pm 0.13$ $2.28$ $0.97$ $0.94$ $0.91$
$1.49\times 10^{-3}$ $1.01\pm 1.20$ $0.98\pm 0.15$ $3.24$ $0.97$ $0.93$ $0.91$
$1.69\times 10^{-3}$ $4\pm 1.24 1.01±0$$1\pm 0.16$ $3.41$ $0.97$ $0.93$ $0.91$